A cyclist is traveling at a fixed speed of 10 miles per hour across the Golden Gate Bridge. She makes it across the bridge in 10 minutes. Based on this scenario, how long is the Golden Gate Bridge? Round the answer to the nearest tenth of a mile. One hour is the same as 60 minute
Answer by mananth(16946)

speed = 10 mph
time to cross distance of bridge= 10 minutes=>1/6 hour
d=r*t
d=10*1/6 = 10/6 miles
=1.7 miles the length of bridge